OpenStudy (mathisfun13):

find the first 4 terms of the geometric sequence a1=40 and r=0.5

OpenStudy (anonymous):

nth_term = a1*r^(term_wanted -1) plug in your numbers

OpenStudy (mathisfun13):

so a1*40?

OpenStudy (anonymous):

1st 40 2nd 40*.5*^1 = 20 3rd 40*.5^2= evaluate

OpenStudy (mathisfun13):

40, 20, 10, 5

OpenStudy (anonymous):

1st is 40 cuzz nth_term = a1*r^(term_wanted -1) 40*.5^( (1)-1 ) 40*.5^0 40*1=40 :note: anynumber^0=1